# In situ measurement of slow solar wind emerging from a pseudostreamer: a conjunction study with Parker Solar Probe and Solar Orbiter

### v2

@@ Abstract
We seek to identify the source regions of the slow solar wind (SSW) through combining models with in situ observations. We leverage an opportune conjunction between Solar Orbiter and Parker Solar Probe (PSP) during PSP Encounter 11 to include compositional diagnostics from the Solar Orbiter heavy ion sensor (HIS) as these variations provide crucial insights into the origin and nature of the solar wind. We use Potential Field Source Surface (PFSS) and Magnetohydrodynamic (MHD) models to connect the observed plasma at PSP and Solar Orbiter to its origin footpoint in the photosphere, and compare these results with the in situ measurements. A very clear signature of a heliospheric current sheet (HCS) crossing as evidenced by enhancements in low FIP elements, ion charge state ratios, proton density, low-Alfvénicity, and polarity estimates validates the combination of modeling, data, and mapping. Fast wind from a small equatorial coronal hole (CH) with low ion charge state ratios, low FIP bias, high-Alfvénicity, and low footpoint brightness mostly fits together, yet includes anomalously low alpha particle abundance. We identify slow wind from many different sources, with broad variation in composition and Alfvénicity. We distinguish between classical non-Alfvénic and high-Alfvénicity SSW and make the association of low-Alfvénicity, decreased alpha-to-proton abundance, high charge state ratios, and FIP bias with streamer wind while intermediate alpha abundance, high-Alfvénicity, and dips in ion charge state ratios correspond to CH boundaries. Through this comprehensive analysis, we highlight the power of multi-instrument conjunction studies in assessing the sources of the solar wind.
@@ 1 Introduction
The large variability observed in the parameters of the slow wind suggests that it originates from a diverse set of source regions, likely influenced by position within the solar cycle, as different regions emerge at various stages. Studies using in situ measurements have found that some slow solar wind shows plasma signatures that are very similar to that of the FSW – high alpha particle abundances (Ohmi et al., 2004), large differential velocities (Stansby et al., 2020a), lower heavy ion charge state ratios and reduced low-FIP enhanced elemental abundances (Stakhiv et al., 2015; D’Amicis et al., 2018), and high Alfvénicity (D’Amicis & Bruno, 2015; Perrone et al., 2020; D’Amicis et al., 2021). These are typical characteristics of the FSW, which means that it is likely that at least some of the slow solar wind originates from CHs, particularly small CHs or the boundaries of larger ones such that the coronal magnetic field strongly over expands. Wang & Sheeley (1990). D’Amicis & Bruno (2015); Stansby et al. (2018, 2020b) and others have shown the existence of two types of slow solar wind: the classical slow solar wind (SSW) with slow speeds and low-Alfvénicity, and the highly Alfvénic slow solar wind (SASW) which shows properties more similar to that of the FSW.
There are a variety of theories using modeling and observational evidence that show the variance in possible sources of the slow solar wind. Eselevich et al. (1999); Riley & Luhmann (2012); Wang et al. (2012); Wang & Panasenco (2019); Panasenco & Velli (2013); Owens et al. (2014); Abbo et al. (2015) and others have shown slow solar wind originating from pseudostreamers – coronal streamers that separate open field lines of the same polarity (Wang et al., 2007). More recent work, such as by D’Amicis et al. (2021); Telloni et al. (2023); Velli et al. (2020); Panasenco et al. (2020), have used PSP and Solar Orbiter data to study the emergence of solar wind from pseudostreamers. Pseudostreamers consist of of multipolar regions of magnetic field which open in a unipolar fashion to interplanetary space, thus separating two coronal holes of the same polarity (Wang et al., 2007). Chitta et al. (2023); Lynch et al. (2023) show that the S-web, a complex web of magnetic separatrices formed by small equatorial CHs reshaping the streamer belt, which reconnects in the middle corona, could drive the slow solar wind. Helmet streamer/streamer belt wind has been shown to contribute to some of the SSW high in the solar atmosphere (5) (Sheeley et al., 1997), which is a different physical process than the steady outflow starting from the photosphere. Antiochos et al. (2011) showed that the heliospheric extension of the streamer belt, the heliospheric current sheet (HCS), is a source of SSW as well. This type of wind from the streamer belt and HCS is generally thought of as the classical SSW, having low Alfvénicity and typical SSW properties. Contrary to the SSW, SASW streams have been associated with origins at CH boundaries or small equatorial CHs (D’Amicis et al., 2021). Additionally, there is the possibility of active region contributions to the SSW during solar maximum (Kasper et al., 2016; Alterman et al., 2018; Alterman & Kasper, 2019; Stansby et al., 2021).
@@
CHs undergo quasi-rigid rotation as the coronal field is continuously undergoing reconnection to remain close to a potential state (Lionello et al., 2005; Wang, 2009). This continuous magnetic reconnection is thought to contribute to to outwardly propagating MHD waves and heating that occurs far above the transition region. In the large loops that contribute to the slow solar wind, MHD waves remain trapped in the transition region, within regions of rapidly diverging open flux where heating is concentrated at low heights (Wang et al., 2009), thus heating the ions leading to enhanced charge state ratios. Additionally, the timescale for fractionation is typically 2 to 3 days, meaning plasma must remain in the corona for this time period to experience fractionation (Laming, 2015). This is why plasma from open field lines that quickly escapes the corona does not see an enhancement in low-FIP abundances, while plasma on closed loops (e.g. those found in active regions forming the streamer belt) does (Geiss et al., 1995; von Steiger et al., 2000). We have discussed above a range of possible slow wind sources thus, we expect to see a wide range of FIP bias in the slow wind from different sources, and relatively low-FIP fractionation in the FSW originating from open magnetic field lines. We use the measurement in conjunction with Alfvénicity, modeling efforts, and other in situ measurements to provide more powerful diagnostics of different types of solar wind and their source region.
In this study, we seek to characterize the source regions of both SASW and SSW through a combination of modeling methods and in situ diagnostics, leveraging a unique conjunction opportunity between inner heliospheric space missions. Understanding the origins of the slow wind will bring us closer to creating a comprehensive model for space weather predictions and allow us to better understand the impact of the solar wind on the near-Earth environment. We begin with a discussion of the spacecraft and remote sensing data that is used in this study in Section 2. We outline the advancements of Parker Solar Probe (PSP; Fox et al. (2016)) and Solar Orbiter (Müller, D. et al. (2020)) that allow for this study and how modeling methods are used in conjunction with in situ data to study the source regions of the solar wind. In Section 3, we discuss the two major magnetic field modeling methods used, potential field source surface (PFSS) and magnetohydrodynamic (MHD) models, and how these methods allow us to study the large-scale coronal magnetic field and identify solar wind source regions. We discuss the observations in Section 4 specifically looking at elemental composition (Section 4.1), particle velocity and abundance ratios (Section 4.2), and modeling efforts (Section 4.3). Lastly, in Section 5, we discuss the main results of the study, then finish with our conclusions and ideas for further work in Section 6.
@@ 2 Data
Our study uses a combination of in situ data from PSP (Fox et al., 2016) and Solar Orbiter (Müller, D. et al., 2020) along with magnetograms from the Helioseismic and Magnetic Imager (HMI; Scherrer et al. (2012)) aboard the Solar Dynamics Observatory (SDO; Pesnell et al. (2012)) and the Global Oscillation Network Group (GONG; Harvey et al. (1996)) during the 11th solar encounter of PSP (E11) in February 18 to March 4, 2022. Through novel measurements, orbital trajectories and instrumentation methods, these instruments have revolutionized our study and understanding of the Sun. The combination of in situ measurements from PSP and Solar Orbiter coupled with modeling results using magnetograms from HMI/SDO and GONG allows us to characterize the solar wind and study its origins. The use of both PSP and Solar Orbiter data is required as Solar Orbiter provides elemental abundance measurements (such as the ratio) and charge state ratios ( and ), which PSP does not, and PSP provides a much better quality calculation of Alfvénicity (see Appendix 8). Additionally, as PSP measures solar wind velocities all below the canonical 500 cutoff during this time period, Solar Orbiter velocity measurements (which are closer to the asymptotic 1AU values) are necessary to characterize the wind as fast or slow according to the canonical 1AU cutoff. Lastly, measuring and identifying features at two substantially different heliocentric distances adds robustness to the analysis by verifying the relevant streams are long-lived and not strongly affected by transients or time evolution.
@@
The PSP mission consists of four instruments to study coronal heating, particle acceleration, and energy flow through the corona: the Electromagnetic Fields Investigation (FIELDS; Bale et al. (2016)), the ‘Solar Wind Electrons, Alphas, and Protons’ (SWEAP; Kasper et al. (2016)) investigation, the Integrated Science Investigation of the Sun (ISIS; McComas et al. (2016)), and the Wide-field Imager for Solar Probe (WISPR; Vourlidas et al. (2016)). From PSP/FIELDS, we use DC magnetic field measurements from the fluxgate magnetometer in radial-tangential-normal (RTN) coordinates (Bale et al., 2016). Proton and alpha particle measurements (density and velocity) from PSP come from the Ion Solar Probe ANalyzer (SPAN-I; Livi et al. (2022)) on PSP/SWEAP. The core electron density was estimated by fitting a bi-Maxwellian distribution along all pitch angles using SPAN-E 3D electron velocity distribution functions (VDFs) (Whittlesey et al., 2020), following a similar method by Romeo et al. (2023). Typically, electron density measurements are calculated from the quasithermal noise spectrum of measurements taken by the Radio Frequency Spectrometer on PSP (Pulupa et al., 2017), however, the instrument was not functioning for a large portion of E11, and thus this measurement is unusable for this time period. The proton and alpha data are cleaned based on the field of view (FOV) of the instrument and through comparison with electron density measurements. All E11 data used in this study is publicly available for download from the PSP/FIELDS 111https://fields.ssl.berkeley.edu/data/ and PSP/SWEAP data archives 222http://sweap.cfa.harvard.edu/Data.html.
@@
The Solar Orbiter mission has a payload of ten in situ and remote sensing instruments. The in situ instruments include an Energetic Particle Detector (EPD; Rodríguez-Pacheco et al. (2020)), Magnetometer (MAG; Horbury et al. (2020)), Radio and Plasma Waves Instrument (RPW; Maksimovic et al. (2021)), and Solar Wind Analyzer (SWA; Owen et al. (2020)). We use Level 2 (science calibrated) DC magnetic field measurements from the magnetometer; and particle and heavy ion measurements from the SWA suite. The DC magnetic field measurements are in RTN coordinates, like the PSP magnetometer measurements. The SWA suite takes high cadence measurements of heavy ion abundances and 3D VDFs of electrons, protons, and alpha particles. The combination of sensors allows for the characterization of the ion and electron bulk properties of the solar wind between 0.28 and 1 AU (Owen et al., 2020). The Heavy Ion Sensor (SWA/HIS; Livi et al. (2023)) determines abundances of heavy ions, which allows us to determine elemental composition and charge state ratios of the plasma, key measurements not available with PSP. Heavy ion measurements such as , , and ratios are a Level 3 data product from SWA/HIS with a resolution time of 10 minutes (Owen et al., 2020; Livi et al., 2023). Using SWA/PAS, we take measured proton fluxes and velocities at a 4-second cadence to characterize the bulk plasma properties of the solar wind. All Solar Orbiter data used in this study can be found on the Solar Orbiter Archive 333https://soar.esac.esa.int/soar/.
@@
In addition to in situ measurements from PSP and Solar Orbiter, we use magnetograms from SDO/HMI and GONG as inputs to our modeling methods (described in Section 3). GONG provides full-disk magnetograms every minute with a resolution of 5 arc seconds, which are publicly available through the NSO website 444https://gong.nso.edu/. HMI provides full-disk Doppler velocity measurements, line of sight photospheric magnetic flux, and continuum images at one arc second resolution with a 45-second cadence along with vector magnetic field measurements at 90 or 135-second cadence (Scherrer et al., 2012). These measurements are produced with the 6177.3nm Fe I line with a resolution of 0.5 arc seconds per pixel, and the solar disk fills almost the entire image frame in HMI data. In addition to HMI, SDO has the Atmospheric Imaging Assembly (AIA; Lemen et al. (2012)), which provides high-resolution full-disk images of the solar corona and transition region at a 12-second resolution. AIA consists of four telescopes taking images at a variety of wavelengths. To produce high-resolution, full-Sun Carrington EUV maps for comparison with models, we use images from the Fe XII (193 Å) bandpass filter aboard SDO/AIA. All SDO/HMI and SDO/AIA data can be accessed through JSOC 555http://jsoc.stanford.edu/.
@@
In order to directly compare measurements between PSP and Solar Orbiter, we plot the observations as a function of projected longitude using ballistic propagation (Figure 1). Snyder & Neugebauer (1966) first proposed this method of connecting an ideal Parker spiral field line from a point in the heliosphere to the corona, assuming a constant solar wind speed. Nolte & Roelof (1973) demonstrated the capabilities of this method to provide a reliable estimate of solar surface footpoints. While this technique does not accurately represent the real solar wind, which has a non-constant speed; Nolte & Roelof (1973); Macneil et al. (2022); Koukras et al. (2022) estimates the error to be less than 10° in longitude as the two corrections, coronal corotation and solar wind acceleration, cancel to first order. We propagate field lines inward from their initial position in the PSP and Solar Orbiter trajectory to their associated latitude and longitude on the source surface at 2.0 using a varying solar wind speed (the in situ measurement) to most accurately represent the real solar wind following Badman et al. (2020) equation (1). This allows for direct comparison of observations as a function of source longitude.
In Figure 1, we show the mapped in situ data and classification of our regions of interest shaded as follows: blue (HCS crossing), pink (FSW), purple (SASW), green (SSW). Wind is classified as fast or slow using the Solar Orbiter velocity measurement with the 500 cutoff (panel (b)), and we identify regions of high (or low) Alfvénicity by looking at the absolute cross helicity () and residual energy () as defined in Equations 1 and 2 (panel (a)). is a proxy for the Alfvénicity of the plasma where higher cross helicity measurements ( values of 1 or -1) are indicative of pure Alfvén waves propagating parallel or antiparallel to the magnetic field. Plasma is considered to have high (or low) Alfvénicity when is above (or below) 0.9, while the residual energy tells us the difference between the energy in velocity and magnetic field fluctuations and which is dominant in the plasma.
@@
We show the full set of the data of interest in Figure 2 as a timeseries of PSP (left column) and Solar Orbiter (right column) measurements: velocity (panels (a) and (f)), scaled proton density (panels (b) and (g)), and scaled radial magnetic field (panels (c) and (h). Both proton density and magnetic field are scaled by for the time period of interest. Panels (d) and (i) show the cross helicity () as calculated for both PSP and Solar Orbiter. We see the measurement from Solar Orbiter does not allow us to characterize the Alfvénicity of the plasma properly, due to a lack of data and the decoherence of the cross helicity measurement with radial distance. Panel (e) shows the Mach number () from PSP while panel (j) shows the SWA/HIS ratio normalized to its photospheric abundance (0.0589) from (Asplund et al., 2021), a strong tracer of the source region. The Alfvénic Mach number the ratio of the bulk solar wind velocity to the Alfvén speed . Periods with below 1 are considered sub-Alfvénic. For this study, we use all the PSP data shown along with the measurements from SWA/HIS on Solar Orbiter and use our ballistic mapping results (Figure 1) to track the mapped features in the timeseries.

### v3

# Compositional metrics of fast and slow Alfvénic solar wind emerging from coronal holes and their boundaries

@@ Abstract
We seek to understand the composition and variability of fast (FSW) and slow Alfvénic solar wind (SASW) emerging from coronal holes (CH). We leverage an opportune conjunction between Solar Orbiter and Parker Solar Probe (PSP) during PSP Encounter 11 to include compositional diagnostics from the Solar Orbiter heavy ion sensor (HIS) as these variations provide crucial insights into the origin and nature of the solar wind. We use Potential Field Source Surface (PFSS) and Magnetohydrodynamic (MHD) models to connect the observed plasma at PSP and Solar Orbiter to its origin footpoint in the photosphere, and compare these results with the in situ measurements. A very clear signature of a heliospheric current sheet (HCS) crossing as evidenced by enhancements in low FIP elements, ion charge state ratios, proton density, low-Alfvénicity, and polarity estimates validates the combination of modeling, data, and mapping. We identify two FSW streams emerging from small equatorial coronal holes (CH) with low ion charge state ratios, low FIP bias, high-Alfvénicity, and low footpoint brightness, yet anomalously low alpha particle abundance for both streams. We identify high-Alfvénicity slow solar wind emerging from the over-expanded boundary of a CH having intermediate alpha abundance, high-Alfvénicity, and dips in ion charge state ratios corresponding to CH boundaries. Through this comprehensive analysis, we highlight the power of multi-instrument conjunction studies in assessing the sources of the solar wind.
@@ 1 Introduction
The large variability observed in the parameters of the slow wind suggests that it originates from a diverse set of source regions, likely influenced by position within the solar cycle, as different regions emerge at various stages. D’Amicis & Bruno (2015); Stansby et al. (2018, 2020b) and others have shown the existence of two types of slow solar wind: the classical slow solar wind (SSW) with slow speeds and low-Alfvénicity, and the highly Alfvénic slow solar wind (SASW) which shows properties more similar to that of the FSW. Studies using in situ measurements have found that some slow solar wind shows plasma signatures that are very similar to that of the FSW – high alpha particle abundances (Ohmi et al., 2004), large differential velocities (Stansby et al., 2020a), lower heavy ion charge state ratios and reduced low-FIP enhanced elemental abundances (Stakhiv et al., 2015; D’Amicis et al., 2018), and high Alfvénicity (D’Amicis & Bruno, 2015; Perrone et al., 2020; D’Amicis et al., 2021). These are typical characteristics of the FSW, which means that it is likely that at least some of the slow solar wind originates from CHs, particularly small CHs or the boundaries of larger ones such that the coronal magnetic field strongly over expands Wang & Sheeley (1990).
@@
CHs undergo quasi-rigid rotation as the coronal field is continuously undergoing reconnection to remain close to a potential state (Lionello et al., 2005; Wang, 2009). This continuous magnetic reconnection is thought to contribute to to outwardly propagating MHD waves and heating that occurs far above the transition region. In the large loops that contribute to the slow solar wind, MHD waves remain trapped in the transition region, within regions of rapidly diverging open flux where heating is concentrated at low heights (Wang et al., 2009), thus heating the ions leading to enhanced charge state ratios. Additionally, the timescale for fractionation is typically 2 to 3 days, meaning plasma must remain in the corona for this time period to experience fractionation (Laming, 2015). This is why plasma from open field lines that quickly escapes the corona does not see an enhancement in low-FIP abundances, while plasma on closed loops (e.g. those found in active regions forming the streamer belt) does (Geiss et al., 1995; von Steiger et al., 2000). We use the measurement in conjunction with Alfvénicity, modeling efforts, and other in situ measurements to provide more powerful diagnostics of different types of solar wind and their source region.
Leveraging a unique conjunction opportunity between inner heliospheric space missions, we use a combination of modeling methods and in situ diagnostics to characterize both FSW and SASW emerging from CHs. We begin with a discussion of the spacecraft and remote sensing data that is used in this study in Section 2. We outline the advancements of Parker Solar Probe (PSP; Fox et al. (2016)) and Solar Orbiter (Müller, D. et al. (2020)) that allow for this study and how modeling methods are used in conjunction with in situ data to study the source regions of the solar wind. In Section 3, we discuss the two major magnetic field modeling methods used, potential field source surface (PFSS) and magnetohydrodynamic (MHD) models, and how these methods allow us to study the large-scale coronal magnetic field and identify solar wind source regions. We discuss the observations in Section 4 specifically looking at elemental composition (Section 4.2), particle velocity and abundance ratios (Section 4.3), and modeling efforts (Section 4.1). Lastly, in Section 5, we discuss the main results of the study, then finish with our conclusions and ideas for further work in Section 6.
@@ 2 Data
Our study uses a combination of in situ data from PSP (Fox et al., 2016) and Solar Orbiter (Müller, D. et al., 2020) along with magnetograms from the Helioseismic and Magnetic Imager (HMI; Scherrer et al. (2012)) aboard the Solar Dynamics Observatory (SDO; Pesnell et al. (2012)) and the Global Oscillation Network Group (GONG; Harvey et al. (1996)) during the 11th solar encounter of PSP (E11) in February 18 to March 4, 2022. Through novel measurements, orbital trajectories and instrumentation methods, these instruments have revolutionized our study and understanding of the Sun. The use of both PSP and Solar Orbiter data is required as Solar Orbiter provides elemental abundance measurements (such as the ratio) and charge state ratios ( and ), which PSP does not, and PSP provides a much better quality calculation of Alfvénicity (see Appendix 8). Additionally, as PSP measures solar wind velocities all below the canonical 500 cutoff during this time period, Solar Orbiter velocity measurements (which are closer to the asymptotic 1AU values) are necessary to characterize the wind as fast or slow according to the canonical 1AU cutoff. Lastly, measuring and identifying features at two substantially different heliocentric distances adds robustness to the analysis by verifying the relevant streams are long-lived and not strongly affected by transients or time evolution.
@@
The PSP mission consists of four instruments to study coronal heating, particle acceleration, and energy flow through the corona. From the Electromagnetic Fields Investigation (FIELDS; Bale et al. (2016)), we use DC magnetic field measurements to understand the global heliospheric field and to validate our models and spacecraft alignments. Proton and alpha particle measurements (density and velocity) from PSP come from the Ion Solar Probe ANalyzer (SPAN-I; Livi et al. (2022)) on the ‘Solar Wind Electrons, Alphas, and Protons’ (SWEAP; Kasper et al. (2016)) investigation. The core electron density was estimated by fitting a bi-Maxwellian distribution along all pitch angles using SPAN-E 3D electron velocity distribution functions (VDFs) (Whittlesey et al., 2020), following a similar method by Romeo et al. (2023). Typically, electron density measurements are calculated from the quasithermal noise spectrum of measurements taken by the Radio Frequency Spectrometer on PSP (Pulupa et al., 2017), however, the instrument was not functioning for a large portion of E11, and thus this measurement is unusable for this time period. The proton and alpha data are cleaned based on the field of view (FOV) of the instrument and through comparison with electron density measurements. All E11 data used in this study is publicly available for download from the PSP/FIELDS 111https://fields.ssl.berkeley.edu/data/ and PSP/SWEAP data archives 222http://sweap.cfa.harvard.edu/Data.html.
@@
The Solar Orbiter mission has a payload of ten in situ and remote sensing instruments. We use Level 2 (science calibrated) DC magnetic field measurements from the Magnetometer (MAG; Horbury et al. (2020)); and particle and heavy ion measurements from the Solar Wind Analyzer (SWA; Owen et al. (2020)) suite. The SWA suite takes high cadence measurements of heavy ion abundances and 3D VDFs of electrons, protons, and alpha particles. The combination of sensors allows for the characterization of the ion and electron bulk properties of the solar wind between 0.28 and 1 AU (Owen et al., 2020). The Heavy Ion Sensor (SWA/HIS; Livi et al. (2023)) determines abundances of heavy ions, which allows us to determine elemental composition and charge state ratios of the plasma, key measurements not available with PSP. Heavy ion measurements such as , , and ratios are a Level 3 data product from SWA/HIS with a resolution time of 10 minutes (Owen et al., 2020; Livi et al., 2023). Using SWA/PAS, we take measured proton fluxes and velocities at a 4-second cadence to characterize the bulk plasma properties of the solar wind. All Solar Orbiter data used in this study can be found on the Solar Orbiter Archive 333https://soar.esac.esa.int/soar/.
@@
We use magnetograms from SDO/HMI and GONG as inputs to our modeling methods (described in Section 3). GONG provides full-disk magnetograms every minute with a resolution of 5 arc seconds, which are publicly available through the NSO website 444https://gong.nso.edu/. HMI provides full-disk Doppler velocity measurements, line of sight photospheric magnetic flux, and continuum images at one arc second resolution with a 45-second cadence along with vector magnetic field measurements at 90 or 135-second cadence (Scherrer et al., 2012). These measurements are produced with the 6177.3nm Fe I line with a resolution of 0.5 arc seconds per pixel, and the solar disk fills almost the entire image frame in HMI data. In addition to HMI, SDO has the Atmospheric Imaging Assembly (AIA; Lemen et al. (2012)), which provides high-resolution full-disk images of the solar corona and transition region at a 12-second resolution. AIA consists of four telescopes taking images at a variety of wavelengths. To produce high-resolution, full-Sun Carrington EUV maps for comparison with models, we use images from the Fe XII (193 Å) bandpass filter aboard SDO/AIA. All SDO/HMI and SDO/AIA data can be accessed through JSOC 555http://jsoc.stanford.edu/.
@@
In order to directly compare measurements between PSP and Solar Orbiter, we plot the observations as a function of projected longitude using ballistic propagation (Figure 1). Snyder & Neugebauer (1966) first proposed this method of connecting an ideal Parker spiral field line from a point in the heliosphere to the corona, assuming a constant solar wind speed. Nolte & Roelof (1973) demonstrated the capabilities of this method to provide a reliable estimate of solar surface footpoints. While this technique does not accurately represent the real solar wind, which has a non-constant speed; Nolte & Roelof (1973); Macneil et al. (2022); Koukras et al. (2022) estimates the error to be less than 10° in longitude as the two corrections, coronal corotation and solar wind acceleration, cancel to first order. We propagate field lines inward from their initial position in the PSP and Solar Orbiter trajectory to their associated latitude and longitude on the source surface at 2.0 using a varying solar wind speed (the in situ measurement) to most accurately represent the real solar wind following Badman et al. (2020) Equation 1. This allows for direct comparison of observations as a function of source longitude.
In Figure 1, we show the mapped in situ data and classification of our regions of interest shaded as follows: blue (HCS crossing), pink (FSW period 1), green (FSW period 2), and purple (SASW). Wind is classified as fast or slow using the Solar Orbiter velocity measurement with the 500 cutoff (panel (b)), and we identify regions of high (or low) Alfvénicity by looking at the absolute cross helicity () and residual energy () as defined in Equations 1 and 2 (panel (a)). is a proxy for the Alfvénicity of the plasma where higher cross helicity measurements ( values of 1 or -1) are indicative of pure Alfvén waves propagating parallel or antiparallel to the magnetic field. Plasma is considered to have high (or low) Alfvénicity when is above (or below) 0.9, while the residual energy tells us the difference between the energy in velocity and magnetic field fluctuations and which is dominant in the plasma.
@@
We show the full set of the data of interest in Figure 2 as a timeseries of PSP (left column) and Solar Orbiter (right column) measurements: velocity (panels (a) and (f)), scaled proton density (panels (b) and (g)), and scaled radial magnetic field (panels (c) and (h). Both proton density and magnetic field are scaled by for the time period of interest. Panels (d) and (i) show the cross helicity () as calculated for both PSP and Solar Orbiter. We see the measurement from Solar Orbiter does not allow us to characterize the Alfvénicity of the plasma properly, due to gaps in the magnetic field observations (panel (h)) which prevent the calculation of and the decoherence of the cross helicity measurement with radial distance. Panel (e) shows the Mach number () from PSP while panel (j) shows the SWA/HIS ratio normalized to its photospheric abundance (0.0589) from (Asplund et al., 2021), a strong tracer of the source region. The Alfvénic Mach number the ratio of the bulk solar wind velocity to the Alfvén speed . Periods with below 1 are considered sub-Alfvénic. For this study, we use all the PSP data shown along with the measurements from SWA/HIS on Solar Orbiter and use our ballistic mapping results (Figure 1) to track the mapped features in the timeseries.